# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v1

@@ 3.2. Container Roles and Implementation Strategies (RQ1)
The next two subsections describe the roles and implementation strategiesS related to containers in multi-cloud environment.
@@ 3.2.1. Container Roles
Table 7 presents a systematic breakdown of the roles of containers in various aspects of multi-cloud environment. It is organized by the “Category Name” and the “Role of Containers” which detail the specific context or usage of containers in that category.
@@
1. Container-based Application Deployment: This category discusses the use of containers for the deployment of various types of applications. This includes web apps, Platform as a Service (PaaS) solutions, data-intensive applications, and Internet of Things (IoT) apps, among others. Containers provide a means of packaging and distributing applications, including their dependencies, across various environments.
2. Container-based Service Hosting: This category refers to the use of containers as a hosting environment for different services. Container-based service hosting allow for scaling and isolation of these services, creating an efficient multi-cloud environment.
3. Container-Based Resource Management: This category discusses the use of containers to manage computational resources. Containers allow for better multi-tenancy solutions, resource sharing, and reduction in complexity, among other benefits.
4. Container Technology Features: This category discusses the specific features and advantages of container technology, such as lightweight virtualization, use in edge computing, and compatibility with orchestration tools like Kubernetes.
5. Containerization in App Development and IoT: The focus of this category is on the use of containers for app development, particularly in the context of IoT. Containers provide a means of encapsulating software units, increasing portability, improving efficiency, and integration with DevOps practices, among other benefits.
6. Performance and Efficiency through Containers: This category discusses how containers can enhance performance and efficiency in a cloud environment. Containers can improve scalability, isolation, and cost-effectiveness, and also aid in platform optimization.
7. Workload Management via Containers: This category is about the use of containers to manage computational workloads across multiple cloud environments. Containers facilitate the migration and execution of these workloads across different platforms.
8. Security and Compliance via Containers: In this category, containers are discussed as a tool for enhancing security and compliance in multi-cloud environment.
9. Other Container Uses: This category includes various other uses of containers, such as data storage, support in architecture, blockchain reliability, scheduling algorithm evaluation, and virtual network composition.
Each of these categories demonstrates a dedicated aspect of how containers can contribute to efficient and effective multi-cloud management. The studies listed in each row (e.g., S01, S02) provide a more in-depth discussion and evidence for the roles and applications of containers.
@@
Container-based Application Deployment Containerized Web-App Deployment S01
Container-Driven App Deployment S05, S08
Containerized PaaS Prototyping S09
Data-Intensive App Deployment via Containers S28
Containerized App Management S29
Container-Based Aneka Service Deployment S51
Container-Driven Microservices Deployment & Scaling S62
Container-Based Multi-Cloud App Deployment & Scaling S74
Containerized IoT App Deployment S85
Container-based Service Hosting Container-Based App Component Hosting S02
Containerized Service Hosting S15
Container-Driven MicroCloud Scaling & Isolation S42
Container-Based Resource Management Container-Based Multi-Tenancy Solution S04
Container-Driven Resource Sharing & Multi-Tenancy S06
Multi-Tenant Service Management via Containers S17
Container-Enabled OS-Level Virtualization S40
Container-Based Complexity Reduction S41
Container Technology Features Container Usage in Edge Computing S22
Container-Enabled Lightweight Virtualization S24
Key Features of Container Technology S25
Container-Based Edge Computing Virtualization S35
Container Leverage in Kubernetes S70
Containerization in App Development & IoT App Development & IoT Integration via Containers S16
Container-Based OpenVAS & Integration of Chef S30
Containers as Standardized Software Units S32
Container Placement & Orchestration in Fog Computing S33
Container Deployment Speed & DevOps Integration S36
Containerizing Apps for Privacy & Compliance S37
Container-Driven Portability & Efficiency Improvement S39
Preventing Vendor Lock-In via Containers S63
Container-Based Extensible App Development S76
Performance & Efficiency via Containers Scalable App Building via Containers S18
Container-Driven SaaS Performance Isolation S23
Cloud Performance Enhancement via Containers S48
Evaluating Performance Impact of Containers S59
Cost Efficiency of Microservices via Containers S68
Container-Based Cloud Platform Optimization S69
Workload Management via Containers Container-Based Virtual Cluster Implementation S21
Multi-Cloud Workload Migration via Containers S60, S65
Seamless Workload Migration via Containers S65
Container-Based Multi-Cloud Platform Execution S79
Other Container Uses Container-Based Azure Data Storage S54
Container & VM Support in Architecture S55
Blockchain-Based Multi-Cloud Reliability via Containers S66
Evaluating Scheduling Algorithm via Containers S67
Container-Based Virtual Network Composition S72
@@ 3.2.2. Implementation Strategies
Table 8 provides an overview of various implementation strategies for container technology in multi-cloud environment. The strategies are grouped into several broad categories, each with a unique focus and varying subcategories. Within these subcategories, each implementation strategy is listed with a concise and revised name for reference.
1. Container Deployment and Management: This category focuses on how containers are deployed and managed in multi-cloud environment. An example from the “Deployment Strategies” subcategory is “Automated Orchestration for Web Applications”, where container deployment is automated for web applications. From the “Container Management” subcategory, the strategy “Container Image Optimization” highlights how the size and efficiency of a container can be optimized.
2. Container Orchestration: In this category, the focus is on coordinating and managing the life cycle of containers in a multi-cloud setup. An instance from the “Cloud Orchestration” subcategory is “SLA-Based Container Strategies”, illustrating how Service Level Agreements (SLAs) can guide the orchestration of containers. In the “Microservices Orchestration” subcategory, “Microservices Architecture in IoT Cloud Apps” represents a way to orchestrate microservices in IoT applications using containers.
3. Container Security: This category focus on the security of containers. An example from the “Container Placement” subcategory is “Secure Placement with Deep Learning”, which employs deep learning techniques to secure the placement of containers. From the “Security Policies” subcategory, the strategy “Access Control and Defense for Container Security” suggests ways to control access and defend containers from security threats.
4. Performance and Scaling: This category addresses the optimization of container performance and the scaling of container-based applications. For example, from the “Performance Optimization” subcategory, “LXC Virtual Cluster for Performance” shows how Linux Containers (LXC) can be used to enhance the performance of virtual clusters. An example from the “Scaling Strategies” subcategory is “Edge Data Streaming and Real-Time Database”, highlighting a strategy for scaling data streaming and real-time databases at the edge.
5. Cloud Service and Networking: This category deals with the management of cloud services and networking in containerized environments. “Multi-Tenant Service with Docker and Cloudant” from the “Cloud Service Management” subcategory illustrates how Docker and Cloudant can be utilized to manage multi-tenant services. From the “Networking” subcategory, the strategy “Edge Scheduling with Location-Awareness” proposes a way to schedule network services at the edge based on location.
6. Data Storage and Persistence: This category discusses strategies related to data management in containerized environments. For instance, the strategy “Firebase Cloud for Real-Time Database” presents an approach that uses Firebase Cloud to manage real-time databases in container-based applications.
@@
Container Deployment and Management Deployment Strategies Automated Orchestration for Web Applications S02
Kubernetes for Containerized Applications S05
Multi-cloud App Development Framework S07
Modular Approach to Container Implementation S08
Deployment with Docker Images S09
Docker Image Creation and Management S31
Container Management Container Image Optimization S32
Efficient Image Reorganization S36
Orchestrating with Docker and Kubernetes S37
Executable Images on Multi-cloud Platforms S79
Container Orchestration Cloud Orchestration SLA-Based Container Strategies S04
Task Division and Scaling in Container Strategy S12
Location-Aware Service Brokering S14
Docker Containers for Multi-tenant Services S17
Resource Acquisition for Aneka Containers S51
Cross-Cloud Container and VM Management S60
Container Orchestration Technologies Overview S33
Microservices Orchestration Microservices Architecture in IoT Cloud Apps S16
ADAPT Deployment of Microservices S27
Microservices Benchmark with Docker and Kubernetes S62
Container Security Container Placement Secure Placement with Deep Learning S06
Container Placement Strategies S39
Security Policies SLA-Oriented Performance Isolation S23
Access Control and Defense for Container Security S25
HIP Integration for Docker Security S77
Secure Connectivity in Multi-tenant Environment S72
Signed URLs for Access Control in Multi-cloud S71
Performance and Scaling Performance Optimization LXC Virtual Cluster for Performance S21
Scalable Multi-layer Architecture with Containers S28
Reactive Auto-scaling and Resource Provisioning S48
Istio for Control Plane Impact and Latency Reduction S59
Scaling Strategies Edge Data Streaming and Real-Time Database S35
MicroCloud Architecture for Efficient Scaling S42
Distributed Workloads with Kubernetes S70
Hybrid Abstractions for Container and Host S40
Cloud Service and Networking Cloud Service Management Multi-Tenant Service with Docker and Cloudant S17
Multi-Cloud Runtime Environment S76
Abstraction Layer for Cloud Service Standardization S63
Networking Edge Scheduling with Location-Awareness S22
Openstack Networking with Kuryr S55
Federated Framework for Cloud Storage S64
Host Identity Protocol for Docker Networking S77
@@
Data Storage and Persistence Data Management Firebase Cloud for Real-Time Database S35
High Availability with SpyStorage MCSS S75
Data Storage with Object Store Service S76
@@ 3.3. Pattern and Strategies (RQ2)
Table 9 presents various patterns and strategies employed in containerized-based applications in multi-cloud environment. These patterns and strategies are divided into various categories and subcategories, each of which highlights different aspects of multi-cloud architecture, management, security, resilience, and migration patterns and strategies. Each category is briefly reported below.
1. Cloud Architectural Patterns and Models: This category reports different architecture patterns, communication and networking patterns, deployment patterns, and service models for containerized-based applications in multi-cloud environment. Architecture Patterns involve several design models for structuring the components of a system, providing a template that guides system designers in building complex systems. The leading patterns identified are microservice architecture and service-oriented architecture (SOA), both of which focus on building applications as suites of independently deployable services. Communication and Networking Patterns relate to how different services within the architecture communicate and network with each other. Service mesh and Service Chaining (SC) are the main patterns observed. Deployment Patterns pertain to different strategies for deploying software systems in multi-cloud environment. The two leading patterns are Blue-Green Deployment and Object Store Service. Service Models involve different ways of providing cloud services, with Platform-as-a-Service (PaaS) and Software-as-a-Servic (SaaS) leading the pack.
2. Cloud Management and Resource Allocation Strategies: This category presents the patterns and strategies related to the management of multi-cloud environment, container management, edge computing, IoT strategies, and resource management. For example, Multi-Cloud Management Strategies include various approaches for managing multiple cloud environments. Multi-Cloud Computing Strategy and Hybrid Cloud Architecture are found leading strategies in this category. Container Management Strategies involve the management of containers in multi-cloud environment, with the Linux Container (LXC) project and Container Engine as the primary patterns identified. Edge Computing and IoT Strategies pertain to the use of edge computing and IoT in multi-cloud environment. Mobile Edge Computing (MEC) and Fog Computing are leading strategies for container management in multi-cloud environment. Resource Management Strategies involve methods for managing resources in the multi-cloud with the help of Kubernetes and Minikube.
3. Cloud Security and Resilience Strategies: In this category, the emphasis is on strategies for ensuring the security and resilience of software systems in multi-cloud environment. Security and Resiliency Patterns are strategies to ensure the system is secure and can recover from failures. The two most prominent patterns are the security-by-design approach and encryption of files. Fault-tolerance Strategies are methods used to ensure that systems can continue to function even when there are failures. Fault-tolerance with Redundant Engines and Multi-cloud Systems Fault-Tolerant Workflow are the leading patterns.
4. Cloud Migration Strategies: This category involves patterns and strategies for moving applications and data between cloud environments and optimizing the migration process. Image-based Migration, for example, is an efficient way to migrate applications because it packages the applications along with their dependencies, configurations, and runtime environment. Migration to a virtualized container, such as Docker, is increasingly popular because it guarantees application portability across various environments and scalability on different cloud platforms.

### v2

@@ 3.2. Container Roles and Implementation Strategies (RQ1)
We report the results related to roles and implementation strategies in the following two subsections.
@@ 3.2.1. Container Roles
We classify the roles of containers in a multi-cloud environment into six categories, as shown in Table 7. Each category is briefly reported below.
1. Container-Based Resource and Service Management: This category includes studies discussing the use of containers for resource management and hosting. The studies underline how the use of containers enables an efficient multi-tenancy solution, sharing of resources, and reduction in complexity due to efficient management of computational resources. It also supports OS-level virtualization and multi-tenant service management, enhancing general resource utilization. In the context of hosting services, containers provide scalable and isolated environments for application components and services. They support microcloud scaling, ensure efficient service delivery, and are vastly used in edge computing for hosting applications closer to end users. Each of these capabilities together shows why containers are very important for optimizing resource management and offering flexible and reliable hosting of services in cloud environments.
@@
2. Container-based Application Deployment: Containers provide a way to package and distribute applications, along with their dependencies, across different environments. This category reports on the use of containers for deploying various types of applications, such as web apps, Platform as a Service (PaaS) solutions, data-intensive applications, and IoT apps.
3. Containerization in App Development and IoT: This category focuses on application development and IoT integration using containers in multi-cloud environment. Key areas include improving deployment speed, enhancing DevOps integration for better portability and efficiency, and enabling standardized software units. Containers are also use to address issues like privacy, compliance, and vendor lock-in. Notable highlights include their role in orchestrating container placement in fog computing and integration with tools like OpenVAS and Chef for enhanced security and automation.
4. Performance and Efficiency through Containers: In this category, we classified studies that discuss how containers improve performance and efficiency. The selected studies highlights their role in enabling scalable application development and enhancing performance isolation for SaaS. Additionally, studies reported that containers contribute to cloud platform optimization, boosting overall cloud performance, and improving cost efficiency specifically while using microservices. Further discussions include evaluating the performance impact of containers across various systems.
5. Container Technology Features: This category also gathers the studies that report the special characteristics and advantages of container technology, to be specific: lightweight virtualization, where the containers improve resource efficiency in edge computing, and applications are brought closer to end users. Additionally, their compatibility with orchestration tools like Kubernetes simplifies container management and deployment across diverse environments.
6. Other Container Uses: This category report several roles of container technology beyond traditional use cases. Key roles include container-based data storage, support for system architecture, and enhancing blockchain reliability in multi-cloud environment. Additionally, containers are used for evaluating scheduling algorithms and enabling virtual network composition, demonstrating their versatility in addressing a wide range of technical challenges.
@@
Container-Based Resource and Service Management Container-Based Multi-Tenancy Solution (Ramos et al., 2021)
Container-Driven Resource Sharing & Multi-Tenancy (Deng et al., 2022)
Multi-Tenant Service Management via Containers (Slominski et al., 2015)
Container-Enabled OS-Level Virtualization (Bélair et al., 2019)
Container-Based Complexity Reduction (Shi et al., 2020)
Containerized Service Hosting (Pham and Pham, 2015)
Multi-Cloud Workload Migration via Containers (Janarthanan et al., 2018; Zeck and Bouroudjian, 2017)
Seamless Workload Migration via Containers (Zeck and Bouroudjian, 2017)
Container-Based Multi-Cloud Platform Execution (Serhiienko and Spillner, 2018)
Container-Driven MicroCloud Scaling & Isolation (Baresi et al., 2016)
Container-based Application Deployment Containerized Web-App Deployment (Almeida et al., 2014)
Container-Driven App Deployment (Gattobigio et al., 2022; Pham et al., 2015)
Containerized PaaS Prototyping (Verdugo et al., 2017)
Data-Intensive App Deployment via Containers (Barna et al., 2017)
Containerized App Management (Gupta et al., 2021)
Container-Based Aneka Service Deployment (Buyya and Barreto, 2015)
Container-Driven Microservices Deployment & Scaling (Rahman and Lama, 2019)
Container-Based Multi-Cloud App Deployment & Scaling (Mfula and Nurminen, 2018)
Containerized IoT App Deployment (Bou Ghantous and Gill, 2021)
Containerization in App Development & IoT App Development & IoT Integration via Containers (Ouyang et al., 2022)
Container-Based OpenVAS & Integration of Chef (Sheridan et al., 2017)
Containers as Standardized Software Units (Al-shammari and Alwan, 2018)
Container Placement & Orchestration in Fog Computing (Sami and Mourad, 2020)
Container Deployment Speed & DevOps Integration (Lorenzo et al., 2019)
Containerizing Apps for Privacy & Compliance (Dimitrijevic et al., 2019)
Container-Driven Portability & Efficiency Improvement (Altran et al., 2022)
Preventing Vendor Lock-In via Containers (Pellegrini et al., 2017)
Container-Based Extensible App Development (Pop et al., 2016)
Performance & Efficiency via Containers Scalable App Building via Containers (Pai and Kumar, 2021)
Container-Driven SaaS Performance Isolation (Wang et al., 2020)
Cloud Performance Enhancement via Containers (Saif et al., 2023)
Evaluating Performance Impact of Containers (Rodigari et al., 2021)
Cost Efficiency of Microservices via Containers (Guerrero et al., 2018a)
Container-Based Cloud Platform Optimization (Kiss, 2018)
Container Technology Features Container Usage in Edge Computing (Wong et al., 2019)
Container-Enabled Lightweight Virtualization (Pahl, 2015)
Key Features of Container Technology (Gao et al., 2017)
Container-Based Edge Computing Virtualization (Carpio et al., 2020)
Container Leverage in Kubernetes (Lazuka et al., 2022)
Other Container Uses Container-Based Azure Data Storage (Bucur et al., 2018)
Container & VM Support in Architecture (Bolivar et al., 2018)
Blockchain-Based Multi-Cloud Reliability via Containers (Aral et al., 2020)
Evaluating Scheduling Algorithm via Containers (Tang, 2021)
Container-Based Virtual Network Composition (Alaluna et al., 2017)
@@ 3.2.2. Implementation Strategies
Container implementation strategies in multi-cloud environment refer to the various approaches and techniques used to deploy, manage, and orchestrate containerized applications across multiple cloud platforms. These strategies ensure efficient, secure, and scalable management of containerized workloads. Table 8 provides an overview of these strategies, categorized into subcategories. Overall, we identified 76 strategies classified into 11 subcategories and 5 categories. Notably, these strategies were identified from only 76 out of 121 selected studies (62.81%). The remaining studies do not explicitly discuss the strategies, or it is unclear what strategies they employed. One possible reason for this could be that many studies focus on high-level discussions of container technology without detailing specific implementation strategies. Additionally, in some cases, the strategies might have been implicitly integrated into broader architectural frameworks, making them difficult to extract or clearly identify.
1. Container Deployment and Management: This category focuses on how containers are deployed and managed in multi-cloud environment. We identified 23 strategies within this category, grouped into three subcategories: Deployment Strategies, Container Lifecycle Management, and Data Management. Deployment strategies, with 10 strategies identified, focus on optimizing the deployment of containers in multi-cloud environment, including approaches such as Kubernetes for containerized applications and Docker image management. Nine strategies related to container lifecycle management address resource monitoring, image optimization, and privacy, with examples like multi-cloud container resource usage monitoring and container image reorganization. The data management subcategory, comprising 4 strategies, focuses on the storage and management of container data in multi-cloud environment, featuring solutions such as Firebase Cloud for real-time databases and blockchain-based container log management. While deployment strategies are well-represented, fewer strategies focus on data management, indicating potential areas for further research in managing data across multi-cloud platforms.
2. Container Orchestration: We identified a total of 20 strategies in this category, classified into two subcategories: Cloud Orchestration and Microservices Orchestration. The Cloud Orchestration subcategory, comprising 13 strategies, focuses on managing containers and resources across multi-cloud environment. These strategies include AI/ML-driven container orchestration, cross-cloud orchestration tools, and Kubernetes CI/CD for multi-cloud setups. The Microservices Orchestration subcategory, consisting of 7 strategies, emphasizes the orchestration of microservices using technologies like Kubernetes and Docker. Key strategies include Kubernetes orchestration for multi-cloud containers and edge computing with Docker and Kubernetes. These approaches are essential for enabling scalable and efficient management of containerized applications, with Cloud Orchestration strategies focusing on adaptive resource management, while Microservices Orchestration supports effective multi-cloud microservices deployment.
3. Container Security: We identified a total of 11 strategies within the Container Security category, classified into two subcategories: Security Policies and Container Placement. The Security Policies subcategory consists of eight strategies and mainly focuses on enhancing container security through mechanisms such as access control, secure connectivity, and deep learning-based policy generation. Key strategies include implementing defense mechanisms, integrating HIP for Docker security, and managing access in multi-cloud environments using signed URLs. These approaches help address security challenges, safeguard containerized applications, and ensure compliance in multi-cloud infrastructures. The Container Placement subcategory consists of three strategies and mainly focuses on optimizing container placement to balance security and efficiency. These strategies include secure placement using deep learning and the dynamic relocation of microservices across clouds. By enabling secure and adaptive placement, these strategies enhance resource utilization, reduce latency, and maintain consistent application performance across cloud environments.
4. Performance Optimization and Scaling Strategies: We identified a total of 12 strategies within the Performance Optimization and Scaling Strategies category, divided into two subcategories: Container Performance Optimization and Scaling Strategies. The Container Performance Optimization subcategory, consisting of seven strategies, focuses on improving container performance in multi-cloud environments. Key approaches include reactive auto-scaling and resource provisioning, AI-driven resource allocation, and multi-cloud resource monitoring. These strategies aim to enhance performance by optimizing resource usage and reducing latency across cloud platforms. The Scaling Strategies subcategory, with five strategies, centers on improving the scalability of containerized applications. Notable approaches include managing distributed workloads with Kubernetes, utilizing hybrid abstractions for container scaling, and implementing MicroCloud architectures for efficient scaling. Together, these strategies support the dynamic adjustment of containerized applications to fluctuating resource demands in multi-cloud setups while maintaining performance and efficiency.
5. Cloud Service and Networking: We identified a total of 10 strategies within the Cloud Service and Networking category, classified into two subcategories: The Cloud Service Integration subcategory, with five strategies, focuses on integrating services across multi-cloud platforms. Key approaches include multi-cloud runtime environments, abstraction layers for cloud service standardization, and the integration of OpenStack and IoT in multi-cloud setups. These strategies aim to simplify and streamline the seamless integration of services across different cloud platforms. The Multi-Cloud Networking Strategies subcategory, also comprising five strategies, addresses networking challenges in multi-cloud environments. Notable strategies include edge scheduling with location awareness, federated frameworks for cloud storage, and TOSCA-based deployment models for multi-cloud setups. These approaches ensure efficient and secure communication between clouds, supporting smooth service operation and data transfer across diverse cloud infrastructures.
@@
Container Deployment and Management Deployment Strategies Automated Orchestration for Web Applications (Jha et al., 2019)
Kubernetes for Containerized APPs (Gattobigio et al., 2022)
Multi-cloud App Development Framework (Zhang et al., 2021)
Modular Approach to Container Implementation (Pham et al., 2015)
Deployment with Docker Images (Verdugo et al., 2017)
Docker Image Creation and Management (Mittal and Risco-Martín, 2017)
Real-time Kubernetes workloads in multi-cloud (Barletta et al., 2024)
AWS multi-cloud container app deployment (Kataru et al., 2023)
Multi-cloud container policy enforcement (Liu and Qin, 2024)
Probes and orchestration for cloud containers (Tundo et al., 2024)
Container Lifecycle Management Container Image Optimization (Al-shammari and Alwan, 2018)
Efficient Image Reorganization (Lorenzo et al., 2019)
Orchestrating with Docker and Kubernetes (Dimitrijevic et al., 2019)
Executable Images on Multi-cloud Platforms (Serhiienko and Spillner, 2018)
Privacy monitoring for containerized clouds (Barati et al., 2023)
Multi-cloud container resource usage monitoring (Horchulhack et al., 2024)
Container ecosystem for multi-cloud deployment (Wong et al., 2023)
Containers in AWS ECS for cloud services (Petrovski and Gusev, 2024)
Container management for multi-cloud MEC (Santos et al., 2023)
IoT container deployment in multi-cloud (Solayman and Qasha, 2023)
@@
Data Management Firebase Cloud for Real-Time Database (Carpio et al., 2020)
High Availability with SpyStorage MCSS (Shen et al., 2017)
Data Storage with Object Store Service (Pop et al., 2016)
Container Orchestration Cloud Orchestration SLA-Based Container Strategies (Ramos et al., 2021)
Task Division and Scaling in Container Strategy (Bisht and Kaur, 2022)
Location-Aware Service Brokering (Chen et al., 2022)
Docker Containers for Multi-tenant Services (Slominski et al., 2015)
Resource Acquisition for Aneka Containers (Buyya and Barreto, 2015)
Cross-Cloud Container and VM Management (Janarthanan et al., 2018)
Container Orchestration Technologies Overview (Sami and Mourad, 2020)
Multi-cloud dynamic resource adaptation (Struhár et al., 2024)
AI/ML orchestration in multi-cloud containers (Sofia et al., 2024)
AI-driven container orchestration across clouds (Benmerar et al., 2023)
Cross-cloud container orchestration tools (Alonso et al., 2023a)
Kubernetes CI/CD for multi-cloud setups (Pandi et al., 2023)
Microservices Orchestration Microservices Architecture in IoT Cloud Apps (Ouyang et al., 2022)
ADAPT Deployment of Microservices (Alonso et al., 2019)
MSA Benchmark with Docker and Kubernetes (Rahman and Lama, 2019)
Kubernetes for power-efficient multi-cloud (Centofanti et al., 2024)
Edge computing with Docker and Kubernetes (Hossain et al., 2023)
Kubernetes for edge-cloud orchestration (Roy and Kalafatis, 2023)
Kubernetes orchestration for multi-cloud containers (Sofia et al., 2023)
Container Security Container Placement Secure Placement with Deep Learning (Deng et al., 2022)
Container Placement Strategies (Altran et al., 2022)
Dynamic microservice relocation across clouds (da Silva et al., 2023)
Security Policies SLA-Oriented Performance Isolation (Wang et al., 2020)
Access Control and Defense for Container Security (Gao et al., 2017)
HIP Integration for Docker Security (Ranjbar et al., 2017)
Secure Connectivity in Multi-tenant Environment (Alaluna et al., 2017)
Signed URLs for Access Control in Multi-cloud (Graupner et al., 2015a)
Performance and Scaling Strategies Performance Optimization LXC Virtual Cluster for Performance (Cui et al., 2016)
Scalable Multi-layer Architecture with Containers (Barna et al., 2017)
Reactive Auto-scaling and Resource Provisioning (Saif et al., 2023)
Istio for Control Plane Impact and Latency Reduction (Rodigari et al., 2021)
Auto-scaling containers across clouds (Dogani et al., 2023)
Multi-cloud resource monitoring and provisioning (Marques et al., 2024)
AI resource allocation in multi-cloud containers (Nagarajan et al., 2023)
Scaling Strategies Edge Data Streaming and Real-Time Database (Carpio et al., 2020)
MicroCloud Architecture for Efficient Scaling (Baresi et al., 2016)
Distributed Workloads with Kubernetes (Lazuka et al., 2022)
Hybrid Abstractions for Container and Host (Bélair et al., 2019)
Cloud Service and Networking Cloud Service Management Multi-Tenant Service with Docker and Cloudant (Slominski et al., 2015)
Abstraction Layer for Cloud Service Standardization (Pellegrini et al., 2017)
OpenStack and IoT in multi-cloud (Merlino et al., 2024)
Data indexing for multi-cloud exposure (Orzechowski et al., 2023)
Networking Edge Scheduling with Location-Awareness (Wong et al., 2019)
Openstack Networking with Kuryr (Bolivar et al., 2018)
Federated Framework for Cloud Storage (Harwalkar et al., 2019a)
Host Identity Protocol for Docker Networking (Ranjbar et al., 2017)
@@ 3.3. Pattern and Strategies (RQ2)
Table 9 presents various patterns and strategies employed in container-based applications in multi-cloud environment. These patterns and strategies are divided into various categories and subcategories, each of which highlights different aspects of multi-cloud architecture, management, security, resilience, and migration patterns and strategies. Notably, several studies (e.g., (Pahl, 2015; Graupner et al., 2015a; Shen et al., 2017; Pham and Pham, 2015)) appear across multiple subcategories because they report different patterns that address distinct aspects of cloud architecture and strategies. Every pattern is classified based on its specific focus, even when originating from the same study. Therefore, we categorized them into separate subcategories to reflect their unique contributions. Each category is briefly reported below.
1. Cloud Architectural Patterns and Models: We identified 48 patterns within this category, grouped into four subcategories: Architecture Patterns, Communication and Networking Patterns, Deployment Patterns, and Service Models. The Architecture Patterns subcategory, with 22 patterns, is the largest, focusing on designing scalable and flexible cloud systems. Leading patterns include Microservice Architecture and Service-Oriented Architecture (SOA), which emphasize modularity and independence of services in multi-cloud environment. Emerging patterns such as AI-Driven Edge-Cloud Architecture and Blockchain-Based Architecture reflect the incorporation of advanced technologies in cloud design. The Communication and Networking Patterns subcategory consists of 11 patterns, with notable strategies such as Service Mesh and Service Chaining (SC) providing essential frameworks for ensuring secure and efficient communication between distributed services across cloud environments. These patterns are critical for enabling seamless integration and operation of cloud services in multi-cloud setups. The Deployment Patterns subcategory, containing 8 patterns, focuses on strategies for deploying applications across multiple cloud platforms. Key patterns include black-Green Deployment, which supports smooth application updates by alternating between two environments, and Object Store Service, which ensures efficient storage of unstructured data across clouds. Advanced orchestration patterns like Decentralized Orchestration Architecture and Policy-Driven Orchestration Architecture further enhance deployment efficiency by automating resource management across clouds. Lastly, the Service Models subcategory comprises 7 models that define various cloud service delivery mechanisms. The leading models, Platform-as-a-Service (PaaS) and Software-as-a-Service (SaaS), provide essential platforms for building, deploying, and managing applications in the cloud without the need for managing underlying infrastructure. Other models, like Function-as-a-Service (FaaS) and Serverless Container-Oriented Architecture, emphasize serverless computing, allowing developers to focus on code rather than infrastructure management.
2. Cloud Management and Resource Allocation Strategies: We identified 30 strategies within this category, grouped into four subcategories: Multi-Cloud Management Strategies, Container Management Strategies, Edge Computing and IoT Strategies, and Resource Management Strategies. The Multi-Cloud Management Strategies subcategory, with 8 strategies, focuses on managing and optimizing resources across multiple cloud environments. Leading strategies include the Multi-cloud Computing Strategy, Hybrid/Multi-cloud Approach, and MAPE-K control loop, all of which emphasize seamless integration and resource optimization across cloud platforms. The Container Management Strategies subcategory contains 7 strategies that address the management of containers in cloud environments. Key strategies include Docker Container Images and Lightweight Virtualization, both of which streamline container management and enhance resource efficiency in multi-cloud setups. The Edge Computing and IoT Strategies subcategory, with 8 strategies, highlights the integration of edge computing and IoT devices with cloud systems. Notable strategies include Mobile Edge Computing (MEC) and Fog Computing, which focus on distributing computation and storage closer to IoT devices for improved performance. Finally, the Resource Management Strategies subcategory includes 7 strategies that focus on efficient resource allocation in cloud environments. Leading approaches like Resource Allocation based on SLA levels, Container Orchestration, Kubernetes, Minikube, and AI-Driven Edge-Cloud Architecture emphasize intelligent resource distribution and orchestration across cloud and edge environments.
3. Cloud Security and Resilience Strategies: We identified 20 strategies within this category, divided into two subcategories: Security and Resiliency Patterns and Fault-tolerance Strategies. The Security and Resiliency Patterns subcategory, comprising 14 strategies, focuses on enhancing security and data resilience in cloud environments. Prominent strategies include Security-by-Design approach, Encryption of files, and Centralized/External ACM Service, which aim to secure data sharing and access control. Emerging patterns like Blockchain-Driven SLA Management Architecture and Hybrid Malware Detection Architecture highlight the use of advanced technologies to ensure resilient cloud infrastructures. The Fault-tolerance Strategies subcategory contains 6 strategies that focus on ensuring system reliability in case of failures. Key strategies include Fault-tolerance with Redundant Engines and Multi-cloud Systems Fault-tolerant Workflow, both of which emphasize maintaining uninterrupted service in multi-cloud environment by building in redundancy and fault-tolerant processes. The RAFT Consensus Algorithm ensures consistency across distributed systems, further improving the resilience of containerized applications in a multi-cloud environment. .
4. Cloud Migration Strategies: We identified 4 strategies within this category, which focus on various approaches to migrating applications and services to the cloud. These strategies include Service-oriented Migration, Application-centric Migration, Image-based Migration, and Migration to a Virtualized Container. Each of these strategies provides a unique method for transferring workloads and data from on-premises or legacy systems to cloud environments. The emphasis is on ensuring a smooth transition with minimal disruption to services, whether through the migration of individual services, entire applications, or the use of virtualized containers. These strategies are essential for organizations looking to leverage cloud technologies while preserving the integrity of their existing systems during the migration process.